# Third-order spectrum of gravitational waves sourced by an excited scalar field

### v1

@@ Abstract
It should be mentioned here that we haven’t considered all two-loop diagrams, which also include cross-correlation terms like $\left\langle h_{-\mathbf{p}}^{(2)}h_{\mathbf{p}}^{(4)}\right\rangle$ and $\left\langle h_{-\mathbf{p}}^{(1)}h_{\mathbf{p}}^{(5)}\right\rangle$. However, using the order parameter we mentioned above, we find that they are of the same order, besides, for higher order diagrams we found that they contain more enhancement factor so they may grow even faster than two loop terms. These facts indicate that perturbative calculations for observable scalar induced GWs during inflation era may be impossible.
Acknowledgements. We thank Xingang Chen, Keisuke Inomata, Jason Kristiano and Atsuhisa Ota for useful comments and discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201501, in part by the National Natural Science Foundation of China under Grant No. 12475067 and No. 12235019.
@@ A.1 Linear order
The Fourier modes of the scalar field perturbations and tensor perturbations are expressed as
@@ A.3 Momentum integrals
As mentioned in the previous chapter, only enhanced k-modes are included in our momentum integrals. Here, I would like to used two examples to illustrate our method. First, the momentum part of $P_{22}$ can be written as
@@
At first glance, a reasonable approximation for a narrow spectrum is
$f\left(|\mathbf{q}-\mathbf{p}|,p\right)=\delta\left[\ln\left(\frac{|\mathbf{q}-\mathbf{p}|}{p_{*}}\right)\right]\delta\left[\ln\left(\frac{p}{p_{*}}\right)\right]f\left(p_{*},p_{*}\right).$ (36)
@@
$\int\frac{\mathrm{d}\theta}{(2\uppi)^{3}}p_{*}^{8}\uppi\frac{\sin^{4}{\theta}}{q}\delta\left(\theta-\theta_{0}\right)=\frac{p_{*}^{4}\uppi}{q}\left(p_{*}^{2}-\frac{q^{2}}{4}\right)^{2}.$ (37)
This result will finally leads us to a $\mathcal{O}(q^{2})$ infrared behaviour which should not happen in local production of GWs. We consider a more physical setting to reveal the problems hidden behind the $\delta$ approximation and to obtain a more reasonable result. We require that only from a specific $p$-mode $p_{*}$ to $p_{*}+d$, the mode functions of $\phi$ are enhanced, thus we have two characteristic momentum scale $q$ and $d$. We consider the integration region $\Omega$ in the momentum space, it is the overlap part of two spherical shell as shown in Fig. 4.
@@
where $k=-2p_{*}\cos{\theta}$. Unfortunately $Y$ goes to infinity as $\theta$ approaches ${\uppi}/{2}$, which means we are facing a infrared divergence and we have to regularize it. To do that we introduce a positive parameter $\theta_{0}$ by hand and integrate it analytically

### v2

@@ Abstract
It should be mentioned here that we haven’t considered all two-loop diagrams, which also include cross-correlation terms like $\left\langle h_{-\mathbf{p}}^{(2)}h_{\mathbf{p}}^{(4)}\right\rangle$ and $\left\langle h_{-\mathbf{p}}^{(1)}h_{\mathbf{p}}^{(5)}\right\rangle$. However, using the order parameter we mentioned above, we find that they are of the same order. Besides, upon examining Feynman diagrams with increasing numbers of loops, we discovered that they incorporate more enhancement factors. Irrespective of the number of loops, there consistently exist terms of order $(2n,4n)$, where $n$ is the number of loops. Thus they may grow even faster than two loop terms. This indicates that in such a case, considering the perturbative expansion of two-loop diagrams or even higher order diagrams will not yield the correct result, and the perturbative method has completely failed here.These facts indicate that perturbative calculations for observable scalar induced GWs during inflation era may be impossible.
Acknowledgements. This work is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201501, in part by the National Natural Science Foundation of China under Grant No. 12075297 and No. 12235019.
@@ A.1 Linear order
The Fourier modes of the scalar field perturbation and tensor perturbation are expressed as
@@ A.3 Momentum integrals
As mentioned in the previous chapter, only enhanced k-modes are included in our momentum integrals. Here, I’d like to used two examples to illustrate our method. First, the momentum part of $P_{22}$ can be written as
@@
At first glance, a reasonable approximation for narrow spectrum is
$f\left(|\mathbf{q}-\mathbf{p}|,p\right)=\delta\left[\ln\left(\frac{|\mathbf{q}-\mathbf{p}|}{p_{*}}\right)\right]\delta\left[\ln\left(\frac{p}{p_{*}}\right)\right]f\left(p_{*},p_{*}\right),$ (36)
@@
$\int\frac{\mathrm{d}\theta}{(2\uppi)^{3}}p_{*}^{8}\uppi\frac{\sin^{4}{\theta}}{q}\delta\left(\theta-\theta_{0}\right)=\frac{p_{*}^{4}\uppi}{q}\left(p_{*}^{2}-\frac{q^{2}}{4}\right)^{2},$ (37)
this result will finally leads us to a $\mathcal{O}(q^{2})$ infrared behaviour which should not happen in local production of GWs. We’d like to consider a more physical setting to reveal the problems hidden behind the $\delta$ approximation and to obtain a more reasonable result. We require that only from a specific $p$-mode $p_{*}$ to $p_{*}+d$, the mode functions of $\phi$ are enhanced, thus we have two characteristic momentum scale $q$ and $d$. We consider the integration region $\Omega$ in the momentum space, it’s the overlap part of two spherical shell as shown in Fig. 4.
@@
where $k=-2p_{*}\cos{\theta}$ ). Unfortunately $Y$ goes to infinity as $\theta$ approaches ${\uppi}/{2}$, which means we are facing a infrared divergence and we have to regularize it. To do that we introduce a positive parameter $\theta_{0}$ by hand and integrate it analytically